OpenStudy (anonymous):

sin 2x-sin 2y/cos 2y=cot(x+y)

OpenStudy (akashdeepdeb):

Is it meant to sin 2x or sin^2x?

OpenStudy (anonymous):

it is sin2x

OpenStudy (akashdeepdeb):

And is the whole term over cos 2y?

OpenStudy (anonymous):

no

OpenStudy (akashdeepdeb):

I don't know! Tag some people!

OpenStudy (anonymous):

i don't get it. do we need to prove that or bring out a relation between x and y?

OpenStudy (anonymous):

prove that

OpenStudy (akashdeepdeb):

@amistre64

OpenStudy (amistre64):

sin 2x - sin 2y/cos 2y = cot(x+y) 2sin(x)cos(x) - 2sin(y)cos(y)/[cos^2(y)-sin^2(y)] = cos(x+y)/sin(x+y) ... = [cos(x)cos(y) - sin(x)sin(y)]/[sin(x)cos(y)+sin(y)cos(x)] thats what i got for a start

OpenStudy (amistre64):

i can never recall the tangent stuff ....

OpenStudy (amistre64):

as is, its not an identity, so its has to do with finding suitable values at best

OpenStudy (anonymous):

kkkkkkk